Last year, Alonzo had
$20,000
to invest. He invested some of it in an account that paid
6%
simple interest per year, and he invested the rest in an account that paid
10%
simple interest per year. After one year, he received a total of
$1280
in interest. How much did he invest in each account?

Answer by addingup(3677)

He had 20,000
Invested x at 6%
Invested 20,000-x at 10%
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
0.06x+0.10(20,000-x) = 1280
0.06x+2,000-0.10x = 1280
-0.04x = -720 divide, and remember -/- = +
x = 18,000 so this is at 6% and
20,000-18,000 = 2,000 is invested at 10%
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
18,000(0.06) = 1,080
2,000(0.10) = 200
_______________-----------
Total. . . . . . 1,280 we have the correct answer
J